# An extension of Steinberg’s Theorem to biquotient pairs of subgroups

### v2

@@
(October 22, 2024)
We study the derived tensor product of the representation rings of subgroups of a given compact Lie group $G$. That is, given two such subgroups $H_{1}$ and $H_{2}$, we study the tensor product of the associated representation rings $\mathrm{R}H_{1}$ and $\mathrm{R}H_{2}$ over the representation ring $\mathrm{R}G$, and prove a vanishing result for the associated higher Tor-groups. This result can be viewed as a natural extension of the Theorem of Steinberg that asserts that the representation rings of maximal rank subgroups of $G$ are free over $\mathrm{R}G$. It my also be viewed as an analogue of a result of Singhof on the cohomology of classifying spaces. We include an immediate application to the complex K-theory of biquotient manifolds.
Consider a compact connected Lie group $G$, with complex representation ring $\mathrm{R}G$. Given two closed connected subgroups $H_{1}$ and $H_{2}$, we can consider their representation rings $\mathrm{R}H_{1}$ and $\mathrm{R}H_{2}$ as $\mathrm{R}G$-modules. The purpose of this note is to prove:
@@
The first motivation is the following result which, following [Panin:twisted] and [Ananyevskiy:homogeneous], we simply refer as Steinberg’s Theorem. [^1] See Remark 17 for a historical discussion.
@@
For any compact connected Lie group $G$ with torsion-free fundamental group, and for any closed connected subgroup $H$ of maximal rank (i. e. $\operatorname{rank}H=\operatorname{rank}G)$, the representation ring $\mathrm{R}H$ is free as an $\mathrm{R}G$-module.
Note that under the stated assumptions, a finitely generated $\mathrm{R}G$-module is free if and only if it is flat (see Lemma 16), so the theorem is equivalent to the assertion that $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H,-)$ vanishes in all positive degrees $i$ for all connected subgroups $H$ of maximal rank. In particular – and this is historically the case of most interest – all higher Tor-groups $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H,\mathbb{Z})$ vanish for connected subgroups $H$ of maximal rank, which immediately implies that a certain spectral sequence of Hodgkin computing the complex K-theory of the homogeneous space $G/H$ collapses (see Section 4 for details). This case corresponds to the special case of Theorem 1 when $H_{1}$ is of maximal rank and $H_{2}$ is trivial, and it is in this sense that Theorem 1 is an extension.
@@
Let $G$ be a compact connected Lie group (with arbitrary fundamental group), with closed connected subgroups $H_{1}$ and $H_{2}$. Consider the rational222 The theorem holds more generally for cohomology with coefficients in any integral domain for which the three cohomology rings $\mathrm{H}^{*}(\mathrm{B}G)$, $\mathrm{H}^{*}(\mathrm{B}H_{1})$ and $\mathrm{H}^{*}(\mathrm{B}H_{2})$ are polynomial algebras with generators in even dimensions. This is always true over $\mathbb{Q}$; see e. g. [McCleary, Theorem 6.38]. cohomology rings $\mathrm{H}^{*}(-):=\mathrm{H}^{*}(-;\mathbb{Q})$ of the classifying spaces $\mathrm{B}G$, $\mathrm{B}H_{1}$ and $\mathrm{B}H_{2}$. If $H_{1}$ intersects every conjugate of $H_{2}$ trivially, then the graded $\operatorname{Tor}$-groups $\operatorname{Tor}_{i}^{\mathrm{H}^{*}(\mathrm{B}G)}(\mathrm{H}^{*}(\mathrm{B}H_{1}),\mathrm{H}^{*}(BH_{2}))$ vanish for all $i>\operatorname{rank}G-(\operatorname{rank}H_{1}+\operatorname{rank}H_{2})$.
We will refer to the condition on $H_{1}$ and $H_{2}$ appearing in Theorem 1 and in the result of Singhof as the strict biquotient condition. This condition arises rather naturally in nature. It is equivalent to the condition that $H_{1}$ acts freely from the left on the homogeneous space $G/H_{2}$ [Singhof:DCM, (1.1)], and hence implies that the double coset space, or biquotient, $H_{1}\backslash G/H_{2}$ is a smooth manifold. We thus refer to $H_{1}\backslash G/H_{2}$ as biquotient manifold.333 In [Singhof:DCM], the strict biquotient condition is referred to as ‘double coset condition’, and biquotient manifolds are called ‘double coset manifolds’. The term ‘biquotient’ used here appears to be vastly more popular.
@@
Singhof’s main application of the above vanishing result was a computation of the cohomology of biquotient manifolds, via an Eilenberg–Moore spectral sequence. Similarly, in Section 4, we will use Theorem 1 to compute the complex K-theory of biquotient manifolds, via the spectral sequence of Hodgkin already mentioned, generalizing the classical computation of the complex K-theory of homogeneous spaces:
@@
Consider a compact connected Lie group $G$ with torsion-free fundamental group, closed connected subgroups $H_{1}$ and $H_{2}$ satisfying the strict biquotient condition, and the associated biquotient manifold $H_{1}\backslash G/H_{2}$. If $\operatorname{rank}H_{1}+\operatorname{rank}H_{2}\geq\operatorname{rank}G-1$, we have isomorphisms
@@
In the maximal rank case, i. e. when $\operatorname{rank}H_{1}+\operatorname{rank}H_{2}=\operatorname{rank}G$, the groups in the second line vanish, i. e. $\mathrm{K}^{1}(G\backslash H_{1}/H_{2})=0$.
It follows, in particular, that under the stated assumptions the natural map from $\mathrm{R}(H_{1}\times H_{2})=\mathrm{R}H_{1}\otimes_{\mathbb{Z}}\mathrm{R}H_{2}$ to $\mathrm{K}^{0}(H_{1}\backslash G/H_{2})$ is surjective. This provides a partial answer to a question raised in [DeVitoGonzalez].
We will establish Theorem 1 in slightly greater generality than stated above.
First, the assumptions on $G$ can be relaxed slightly. Let’s say a Lie group $G$ is good if it is compact and connected and if $\mathrm{R}G$ is a polynomial ring tensored with a Laurent ring. All compact connected Lie groups with torsion-free fundamental group are good (Lemma 14), but not conversely. Steinberg established the theorem cited above under this more general assumption, see Theorem 13 below, and our Theorem 1 likewise holds for all good $G$.
Second, the strict biquotient condition can be relaxed as follows. Two subgroups $H_{1}$ and $H_{2}$ of $G$ satisfy the lax biquotient condition if every intersection of $H_{1}$ with a conjugate of $H_{2}$ is central in $G$. In this case, the action of $H_{1}$ on $G/H_{2}$ is still effectively free, and the double coset space $H_{1}\backslash G/H_{2}$ is still a smooth manifold. Theorem 1 also holds in this situation, but with an obvious correction term. Given any pair of subgroups $H_{1}$, $H_{2}$, we define the intersection rank $\mathrm{interrank}(H_{1},H_{2})$ as the maximum of the ranks of the intersections $H_{1}\cap gH_{2}g^{-1}$ as $g$ ranges over all elements of $G$. With this definition in place, the result reads as follows:
@@
We say a few words about the proof. There is little relation of our proof with Singhof’s proof, which makes crucial use of the grading in cohomology.
It turns out that the main difficulty is in establishing Theorem 3 in the maximal rank case, i. e. when $\operatorname{rank}G=\operatorname{rank}H_{1}+\operatorname{rank}H_{2}-\mathrm{interrank}(H_{1},H_{2})$. When $G$ is a torus, this case can be verified by a direct calculation using the spectral sequence of a double complex. The reader may easily reconstruct this calculation by stripping away all unnecessary complications from the proof of Proposition 32 below. It is also easy, for general $G$, to reduce to the case that $H_{1}$ and $H_{2}$ are tori; see Proposition 31. However, we have found no easy argument for passing from the case that the ambient group $G$ is a torus to the more general case. Instead, we try to salvage as much as possible from the calculation that is possible when $G$ is a torus. Two important ingredients are: (1) Segal’s theory of supports of prime ideals in representation rings [Segal:RG], which allows us to translate the biquotient condition into commutative algebra (Proposition 22), and (2) a “reduction to the diagonal”, which allows us to assume that $\mathrm{R}H_{2}$ is a quotient of $\mathrm{R}G$ by a regular sequence (see Proposition 25 and the proof of Proposition 33).
The passage from the maximal rank case to the general case is achieved by enlarging $H_{1}$ (see Proposition 29 and the proof of Proposition 34). However, in general this enlargment process destroys the biquotient condition (Remark 30). This forces us to prove Theorem 3 in somewhat greater generality than stated above, see Proposition 34. In short, we show that the Tor-groups always vanish at certain primes of $\mathrm{R}G$, and for subgroups satisfying the biquotient condition, these are the only primes that matter.
@@ Acknowledgements
I am very grateful to Wilhelm Singhof for asking me whether Theorem 1 holds, and to David González-Álvaro and Panagiotis Konstantis for repeatedly motivating me to come up with an answer. Jeff Carlson and Oliver Goertsches shared their thoughts on various preliminary approaches to the problem. The special case of Theorem 1 in which $\operatorname{rank}H_{1}=1$ was established by Alessandra Wiechers in her Master’s thesis [Wiechers], already using Segal’s theory of supports. Leopold Zoller outlined a strategy to me by which the first isomorphism of Corollary 2 might alternatively be established in the case when $G$ is a product of special unitary and symplectic groups, and $H$ is of maximal rank, starting from the cohomological result of Singhof. Jakob Bergqvist shared some helpful insights regarding Lemma 11. Our use of the “reduction to the diagonal” is inspired by [Serre:LA, Chapter V.C, § 4].
@@ 1.1 Abelian groups
The first functor, which takes a compact abelian group $S$ to its character group ${\operatorname{Hom}(S,S^{1})}$, is a contravariant equivalence by Pontryagin duality. The second functor takes an abelian group $A$ to the associated group ring $\mathbb{Z}[A]$. In other words, $\mathrm{R}S\cong\mathbb{Z}[{\operatorname{Hom}(S,S^{1})}]$ for any compact Abelian group $S$. For example, when $T$ is a torus of rank $r$, i. e. when $T\cong(S^{1})^{r}$, we have $\operatorname{Hom}(T,S^{1})\cong\mathbb{Z}^{r}$. So $\mathrm{R}(T)$ is isomorphic to $\mathbb{Z}[x_{1}^{\pm 1},\dots,x_{r}^{\pm 1}]$, a Laurent ring on $r$ generators. We will need the following consequences of this explicit description.
@@
    Any subtorus $S\subseteq T$ is a direct factor. For such a subtorus $S$, the kernel of the restriction $\mathrm{R}T\to\mathrm{R}S$ is generated by a regular sequence of length $\operatorname{rank}(T)-\operatorname{rank}(S)$.
@@
    Under this correspondence, $\operatorname{rank}(K_{S})=r-\operatorname{rank}(S)$, the subtori of $T$ correspond to the direct summands of $\mathbb{Z}^{r}$, and $K_{S_{1}\cap S_{2}}=K_{S_{1}}+K_{S_{2}}$.
@@ 1.2 Arbitrary compact groups
The first observation follows directly from the definition of $\mathrm{R}G$ and is only mentioned here because we will implicitly use it when applying Theorem 20 later. For (b), see [Segal:RG, Corollary 3.3].
For (c), consider first a compact abelian group $S$. Pullback along the projection to the identity component $S\twoheadrightarrow S_{0}$ defines a finite ring extension $\mathrm{R}S_{0}\hookrightarrow\mathrm{R}S$. So $\dim(\mathrm{R}S)=\dim(\mathrm{R}S_{0})$. As $\mathrm{R}S_{0}=\mathbb{Z}[x_{1}^{\pm 1},\dots,x_{r}^{\pm 1}]$ with $r=\operatorname{rank}S_{0}=\operatorname{rank}S$, the claim follows in this case.
For any compact connected $G$, the restriction to a maximal torus $T$ defines an integral extension $\mathrm{R}G\to\mathrm{R}T$ (c. f. Lemma 8 below). So the result for $G$ follows from the result for $T$ which we have just established. For general compact $G$, we similarly have an integral extension $\mathrm{R}G\to\prod_{S}\mathrm{R}S$, where the product is over a finite number of compact abelian groups $S$ [Segal:RG, after Corollary (3.3)]. So $\dim\mathrm{R}G=\max_{S}\dim\mathrm{R}S$. More precisely, the groups $S$ range over all conjugacy classes of what Segal calls “Cartan subgroups” of $G$ (though note thas this terminology conflicts with modern usage). Not every such Cartan subgroup is of maximal rank, as the example in the final remark of [Segal:RG, § 1] shows. However, any maximal torus of the identity component of $G$ is a Cartan subgroup, so at least one of the groups $S$ has maximal rank, and the claim follows. ∎

### v3

@@
(23rd January 2026)
We study the derived tensor product of the representation rings of subgroups of a given compact Lie group $G$. That is, given two such subgroups $H_{1}$ and $H_{2}$, we study the tensor product of the associated representation rings $\mathrm{R}H_{1}$ and $\mathrm{R}H_{2}$ over the representation ring $\mathrm{R}G$, and prove a vanishing result for the associated higher Tor-groups. This result can be viewed as a natural extension of the Theorem of Steinberg that asserts that the representation rings of maximal rank subgroups of $G$ are free over $\mathrm{R}G$. It may also be viewed as an analogue of a result of Singhof on the cohomology of classifying spaces. We include an immediate application to the complex K-theory of biquotient manifolds.
Consider a compact connected Lie group $G$, with complex representation ring $\mathrm{R}G$. Given two closed connected subgroups $H_{1}$ and $H_{2}$, we can consider their representation rings $\mathrm{R}H_{1}$ and $\mathrm{R}H_{2}$ as $\mathrm{R}G$-modules. The purpose of this paper is to prove:
@@
The first motivation is the following result which, following [Panin:twisted] and [Ananyevskiy:homogeneous], we simply refer to as Steinberg’s Theorem. [^1] See Remark 18 for a historical discussion.
@@
For any compact connected Lie group $G$ with torsion-free fundamental group, and for any closed connected subgroup $H$ of maximal rank (i. e., $\operatorname{rank}H=\operatorname{rank}G)$, the representation ring $\mathrm{R}H$ is free as an $\mathrm{R}G$-module.
Note that under the stated assumptions, a finitely generated $\mathrm{R}G$-module is free if and only if it is flat (see Lemma˜17), so Steinberg’s Theorem is equivalent to the assertion that $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H,-)$ vanishes in all positive degrees $i$ for all connected subgroups $H$ of maximal rank. In particular – and this is historically the case of most interest – all higher Tor-groups $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H,\mathbb{Z})$ vanish for connected subgroups $H$ of maximal rank, which immediately implies that a certain spectral sequence of Hodgkin computing the complex K-theory of the homogeneous space $G/H$ collapses (see Section˜4 for details). This case corresponds to the special case of Theorem˜1 when $H_{1}$ is of maximal rank and $H_{2}$ is trivial, and it is in this sense that Theorem˜1 is an extension.
@@
Let $G$ be a compact connected Lie group (with arbitrary fundamental group), with closed connected subgroups $H_{1}$ and $H_{2}$. Consider the rational [^2] The theorem holds more generally for cohomology with coefficients in any integral domain for which the three cohomology rings $\mathrm{H}^{*}(\mathrm{B}G)$, $\mathrm{H}^{*}(\mathrm{B}H_{1})$ and $\mathrm{H}^{*}(\mathrm{B}H_{2})$ are polynomial algebras with generators in even dimensions. This is always true over $\mathbb{Q}$; see e. g. [McCleary, Theorem 6.38]. cohomology rings $\mathrm{H}^{*}(-):=\mathrm{H}^{*}(-;\mathbb{Q})$ of the classifying spaces $\mathrm{B}G$, $\mathrm{B}H_{1}$ and $\mathrm{B}H_{2}$. If $H_{1}$ intersects every conjugate of $H_{2}$ trivially, then the graded $\operatorname{Tor}$-groups $\operatorname{Tor}_{i}^{\mathrm{H}^{*}(\mathrm{B}G)}(\mathrm{H}^{*}(\mathrm{B}H_{1}),\mathrm{H}^{*}(BH_{2}))$ vanish for all $i>\operatorname{rank}G-(\operatorname{rank}H_{1}+\operatorname{rank}H_{2})$.
We will refer to the condition on $H_{1}$ and $H_{2}$ appearing in Theorem˜1 and in the result of Singhof as the strict biquotient condition. This condition arises rather naturally. It is equivalent to the condition that $H_{1}$ acts freely from the left on the homogeneous space $G/H_{2}$ [Singhof:DCM, (1.1)], and hence implies that the double coset space, or biquotient, $H_{1}\backslash G/H_{2}$ is a smooth manifold. We thus refer to $H_{1}\backslash G/H_{2}$ as biquotient manifold. [^3] In [Singhof:DCM], the strict biquotient condition is referred to as ‘double coset condition’, and biquotient manifolds are called ‘double coset manifolds’. The term ‘biquotient’ used here appears to be vastly more popular.
@@
Singhof’s main application of the above vanishing result was a computation of the cohomology of biquotient manifolds, via an Eilenberg–Moore spectral sequence. Similarly, in Section˜4, we will use Theorem˜1 to compute the complex K-theory of biquotient manifolds, via the spectral sequence of Hodgkin already mentioned, generalizing the classical computation of the complex K-theory of homogeneous spaces:
@@
Consider a compact connected Lie group $G$ with torsion-free fundamental group, closed connected subgroups $H_{1}$ and $H_{2}$ satisfying the strict biquotient condition, and the associated biquotient manifold $H_{1}\backslash G/H_{2}$. If $\operatorname{rank}H_{1}+\operatorname{rank}H_{2}\geq\operatorname{rank}G-1$, we have the following isomorphisms:
@@
In the maximal rank case, i. e., when $\operatorname{rank}H_{1}+\operatorname{rank}H_{2}=\operatorname{rank}G$, the groups in the second line vanish, i. e., $\mathrm{K}^{1}(H_{1}\backslash G/H_{2})=0$.
Note that all concepts under consideration here depend only on the conjugacy classes of $H_{1}$ and $H_{2}$: The representation rings of conjugate subgroups $H_{i}$ and $H_{i}^{\prime}$ of $G$ are isomorphic as $\mathrm{R}G$-modules (see Remark˜8), and the biquotients $H_{1}\backslash G/H_{2}$ and $H_{1}^{\prime}\backslash G/H_{2}^{\prime}$ are diffeomorphic. Corollary˜2 implies that, under the stated assumptions, the natural map from $\mathrm{R}(H_{1}\times H_{2})\cong\mathrm{R}H_{1}\otimes_{\mathbb{Z}}\mathrm{R}H_{2}$ to $\mathrm{K}^{0}(H_{1}\backslash G/H_{2})$ is surjective. This provides a partial answer to a question raised in [DeVitoGonzalez].
We will establish Theorem˜1 in slightly greater generality than stated above.
First, the assumptions on $G$ can be relaxed slightly. Let’s say a Lie group $G$ is good if it is compact and connected and if $\mathrm{R}G$ is a polynomial ring tensored with a Laurent ring. All compact connected Lie groups with torsion-free fundamental group are good (Lemma˜15), but not conversely. Steinberg established the theorem cited above under this more general assumption (see Theorem˜14 below), and our Theorem˜1 likewise holds for all good $G$.
Second, the strict biquotient condition can be relaxed as follows. Two subgroups $H_{1}$ and $H_{2}$ of $G$ satisfy the lax biquotient condition if every intersection of $H_{1}$ with a conjugate of $H_{2}$ is central in $G$. In this case, the action of $H_{1}$ on $G/H_{2}$ is still effectively free, and the double coset space $H_{1}\backslash G/H_{2}$ is still a smooth manifold. Theorem˜1 also holds in this situation, but with an obvious correction term. Given any pair of subgroups $H_{1}$, $H_{2}$, we define the intersection rank $\mathrm{interrank}(H_{1},H_{2})$ as the maximum of the ranks of the intersections $H_{1}\cap gH_{2}g^{-1}$ as $g$ ranges over all elements of $G$. With this definition in place, the result reads as follows:
@@
We say a few words about the proof. Despite the close analogy of the statements, there is very little relation of our proof of Theorem˜3 with Singhof’s proof of the result cited above, which makes crucial use of the grading in cohomology.
We first establish Theorem˜3 in the maximal rank case, i. e., when $\operatorname{rank}G=\operatorname{rank}H_{1}+\operatorname{rank}H_{2}-\mathrm{interrank}(H_{1},H_{2})$. When the ambient group $G$ is a torus, this case can be verified by a direct calculation using the spectral sequence of a double complex. The reader may easily reconstruct this calculation by stripping away all unnecessary complications from the proof of Proposition˜35 below. It is also easy, for general $G$, to reduce to the case that $H_{1}$ and $H_{2}$ are tori; see Proposition˜34. However, we have found no simple argument for passing from the statement for ambient tori to the statement for more general ambient groups $G$. Instead, we try to salvage as much as possible from the actual calculation for tori, and generalize the whole argument. One key ingredient is a “reduction to the diagonal”, which allows us to assume that $\mathrm{R}H_{2}$ is a quotient of $\mathrm{R}G$ by a regular sequence (see Proposition˜27 and the proof of Proposition˜36).
The passage from the maximal rank case to the general case is achieved by enlarging $H_{1}$ (see Proposition˜31 and the proof of Proposition˜37). However, in general, this enlargment process destroys the biquotient condition (Remark˜33). This forces us to prove Theorem˜3 in somewhat greater generality than stated above: we show that, for arbitrary pairs of connected subgroups, the relevant Tor groups vanish when localized at primes supporting the representation ring $\mathrm{R}(\mathrm{Z}G)$ of the centre of $G$ (see Proposition˜37). The key for extracting Theorem˜3 from this result is Segal’s theory of supports of prime ideals in representation rings [Segal:RG], which allows us to translate the biquotient condition into commutative algebra (Proposition˜23): the lax biquotient condition implies that the only primes of $\mathrm{R}G$ supporting both $\mathrm{R}H_{1}$ and $\mathrm{R}H_{2}$ are those primes also supporting $\mathrm{R}(\mathrm{Z}G)$ (Example˜25). Thus, we may deduce that the relevant Tor groups vanish by arguing one prime $\mathfrak{p}$ at a time: if $\mathfrak{p}$ supports $\mathrm{R}(\mathrm{Z}G)$, the Tor groups vanish at $\mathfrak{p}$ by the more general result; if not, they vanish because at least one of $\mathrm{R}H_{1}$ and $\mathrm{R}H_{2}$ do so.
@@ Acknowledgements
I am very grateful to Wilhelm Singhof for asking me whether Theorem˜1 holds, and to David González-Álvaro and Panagiotis Konstantis for repeatedly motivating me to come up with an answer. Jeff Carlson and Oliver Goertsches shared their thoughts on various preliminary approaches to the problem. The special case of Theorem˜1 in which $\operatorname{rank}H_{1}=1$ was established by Alessandra Wiechers in her Master’s thesis [Wiechers], already using Segal’s theory of supports. Leopold Zoller outlined a strategy to me by which the first isomorphism of Corollary˜2 might alternatively be established in the case when $G$ is a product of special unitary and symplectic groups, and $H$ is of maximal rank, starting from the cohomological result of Singhof. Jakob Bergqvist shared some helpful insights regarding Lemma˜12. Our use of the “reduction to the diagonal” is inspired by [Serre:LA, Chapter V.C, § 4]. Finally, I would like to thank the anonymous referee for careful reading and meticulous reports that have led to numerous improvements.
@@ 1.1 Abelian groups
The first functor, which takes a compact abelian Lie group $S$ to its character group $\operatorname{Hom}(S,S^{1})$}, is a contravariant equivalence by Pontryagin duality. The second functor takes an abelian group $A$ to the associated group ring $\mathbb{Z}[A]$. In other words, $\mathrm{R}S\cong\mathbb{Z}[{\operatorname{Hom}(S,S^{1})}]$ for any compact abelian Lie group $S$. For example, when $T$ is a torus of rank $r$, i. e., when $T\cong(S^{1})^{r}$, we have $\operatorname{Hom}(T,S^{1})\cong\mathbb{Z}^{r}$. So $\mathrm{R}T$ is isomorphic to $\mathbb{Z}[x_{1}^{\pm 1},\dots,x_{r}^{\pm 1}]$, a Laurent ring on $r$ generators. We will need the following consequences of this explicit description.
@@
    Any subtorus $S\subseteq T$ is a direct factor. For such a subtorus $S$, the kernel of the restriction $\mathrm{R}T\to\mathrm{R}S$ is generated by a regular sequence of length $\operatorname{rank}{T}-\operatorname{rank}{S}$.
@@
    Under this correspondence, $\operatorname{rank}{K_{S}}=r-\operatorname{rank}{S}$, the subtori of $T$ correspond to the direct summands of $\mathbb{Z}^{r}$, and $K_{S_{1}\cap S_{2}}=K_{S_{1}}+K_{S_{2}}$.
@@ 1.2 Arbitrary compact groups
The first observation follows directly from the definition of $\mathrm{R}G$ and is only mentioned here because we will implicitly use it when applying Theorem˜21 later. For (b), see [Segal:RG, Corollary 3.3].
For (c), consider first a compact abelian group $S$. Pullback along the projection to the identity component $S\twoheadrightarrow S_{0}$ defines a finite ring extension $\mathrm{R}S_{0}\hookrightarrow\mathrm{R}S$. So $\dim{\mathrm{R}S}=\dim{\mathrm{R}S_{0}}$. As $\mathrm{R}S_{0}=\mathbb{Z}[x_{1}^{\pm 1},\dots,x_{r}^{\pm 1}]$ with $r=\operatorname{rank}S_{0}=\operatorname{rank}S$, the claim follows in this case.
For any compact connected $G$, the restriction to a maximal torus $T$ defines an integral extension $\mathrm{R}G\to\mathrm{R}T$ (see Lemma˜9 below). So the result for $G$ follows from the result for $T$ which we have just established. For general compact $G$, we similarly have an integral extension $\mathrm{R}G\to\prod_{S}\mathrm{R}S$, where the product is over a finite number of compact abelian groups $S$ [Segal:RG, proof of Proposition (3.2)]. So $\dim\mathrm{R}G=\max_{S}\dim\mathrm{R}S$. More precisely, the groups $S$ range over all conjugacy classes of what Segal calls “Cartan subgroups” of $G$ (though note that this terminology conflicts with modern usage). Not every such Cartan subgroup is of maximal rank, as the example in the final remark of [Segal:RG, § 1] shows. However, any maximal torus of the identity component of $G$ is a Cartan subgroup, so at least one of the groups $S$ has maximal rank, and the claim follows. ∎
@@
Remark 8.
The representation ring $\mathrm{R}H^{\prime}$ of a subgroup $H^{\prime}$ conjugate to $H$ is isomorphic to $\mathrm{R}H$ as an $\mathrm{R}G$-module. This follows from the fact that conjugation by an inner automorphism induces the identity on $\mathrm{R}G$.